# AuscultaBase: A Foundational Step Towards AI-Powered Body Sound Diagnostics

### v1

@@ 4 AuscultaBase-Corpus: Pretraining Datasets
Table 1: Statistics of the pretraining datasets (SR: sampling rate; Duration: mean [min, max]), where the sound type is categorized into lung (L), heart (H), and bowel (B) sounds.
5 AuscultaBase-Model: Auscultation Foundation Model
We utilize the aforementioned combination of datasets in Table 1 to pre-train our auscultation model. For the datasets that have been split into training and validation sets, we reserve the validation ones as our validation sets. For others, we shuffle the samples and randomly select 10% for validation.
As the result of the nature of body sounds, the recordings of them usually share similar patterns. Therefore, to make the model distinguish the nuances among recordings and capture the abnormality of each audio, we adopt the contrastive learning-based objective to pre-train our model. Specifically, two segments from the same recording are regarded as positive pairs, otherwise negative pairs. The rationale behind this choice is that if an encoder can capture nuances among different sources, it is expected to encode recognizable and generalizable acoustic features for downstream evaluation.
To adapt to the varied length of recordings as exhibited in Table 1, we refer to Chen et al. (2022) and utilize a transformer-based network as the encoder to extract features. Then a projector is adopted to map the features into a lower dimension, and the bilinear similarity will be calculated among obtained low-dimensional representations. The optimization objective aims to maximize the similarity between positive pairs and minimize it for negative pairs. Specifically, due to inherent variations in audio length within an individual batch, we employ random cropping to obtain segments for training. To guarantee the difficulty of the optimization target, we construct a separate dataloader for each dataset and ensure the audio clips from each batch are from the same dataset. The batches are sampled from different data loaders randomly.
The cropped length is set as 320 ms for Korean and Bowel Sound while other datasets share the length of 640 ms. We project the encoded representations into a 512-dimensional space and the model for 200 epochs. We employ an Adam optimizer with a learning rate of 1e-4 which decayed by 0.99 every one epoch.
6 AuscultaBase-Bench: Auscultation Benchmark
To assess the auscultation capability of existing auscultation foundation models, we construct AuscultaBase-Bench, a novel benchmark on all the datasets we have collected as shown in Table 1. It is noted that we have previously excluded all the test sets in the pertaining state to avoid data leakage.
@@ 6 AuscultaBase-Bench: Auscultation Benchmark
ID Dataset Sound Type Task Type Class Name Data Distribution
T1 SPRSound L MC Normal / Adventitious / Poor Quality 2324 / 1000 / 230
T2 SPRSound L MC Normal / CAS / DAS / CAS&DAS / Poor Quality 2324 / 368 / 480 / 152 / 230
T3 HF Lung L BC Normal / Abnormal 52444 / 29489
T4 HF Lung L MC Inhalation / Exhalation / CAS / DAS 34095 / 18349 / 13883 / 15606
T5 HF Lung L MC Inhalation / Exhalation / Wheeze / Stridor / Rhonchi / Crackle 34095 / 18349 / 8457 / 686 / 4740 / 15606
T6 ICBHI 2017 L MC Normal / Crackle / Wheeze / Crackle&Wheeze 3642 / 1864 / 886 / 506
T7 Lung Sound L MC Normal / Crepitation / Wheeze / Crackle / Bronchi / Wheeze&Crackle / Bronchi&Crackle 105 / 69 / 123 / 24 / 3 / 6 / 6
T8 Circor 2022 H MC Present / Absent / Unknown 363 / 2391 / 156
T9 ICBHI 2017 L MC Healthy / Bronchiectasis / Bronchiolitis / COPD / Pneumonia / URTI 35 / 16 / 13 / 793 / 37 / 23
@@
T10 Lung Sound L ML Normal / Asthma / Pneumonia / COPD / BRON / Heart failure / Lung fibrosis / Pleural effusion 105 / 99 / 15 / 33 / 9 / 63 / 18 / 6
T11 Respiratory Database@TR L MC COPD0 / COPD1 / COPD2 / COPD3 / COPD4 72 / 60 / 84 / 84 / 204
@@
T12 Korean H MC Normal / Aortic Stenosis / Mitral Regurgitation / Mitral Stenosis / Murmur in Systole 200 / 200 / 200 / 200 / 200
T13 Cinc 2016 H BC Normal / Abnormal 2575 / 665
T14 Circor 2022 H BC Normal / Abnormal 1632 / 1531
T15 HSDReport H BC Normal / Abnormal 247 / 2028
T16 Bowel Sound B R - 0$\sim$43
@@
Table 2: Downstream task characteristics grouped by function category (T1-T8: abnormality detection. T9-T15: disease diagnosis, T16: activity detection). The task type is categorized into binary classification (BC), multi-class classification (MC), multi-label classification (ML), and regression (R) tasks. Data Distribution represents the numerical range of the frequency of bowel sounds for T16, while for T1-T15 it denotes the number of samples corresponding to the class names.
6.1 Tasks
@@
Our benchmark consists of a total of 16 tasks involving binary classification, multi-class classification, multi-label classification, and regression, as demonstrated in Table 2. We assort these tasks by their function as abnormality detection, disease diagnosis, and activity detection tasks.
@@
tasks (T1-T8) aim to distinguish whether the body sound is normal and which normality or abnormality it belongs to. For example, T8 targets to identify if any murmur exists in the heart sound, while T4 requires the model to identify whether the patient is inhaling or exhaling and whether the adventitious sounds are continuous (CAS) or discrete (DAS) if they exist. The abnormality of body sounds may indicate unusual health conditions but still require clinicians to confirm and report the diagnosis.
@@
tasks (T9-T15) focus on distinguishing if the patient is healthy and which disease the patient may suffer from. For example, T12 aims to distinguish whether the heart sound is healthy or reveals any symptoms of cardiac diseases. T11 requires models to identify the level of chronic obstructive pulmonary disease (COPD) which ranges from 0 to 4. It is noted that we exclude “LRTI” and “Asthma” from the original dataset for T9 because The sample proportions for these two categories are rather small and do not appear in the test set. These tasks require the auscultation model to integrate the entire recording to analyze if any diagnosis exists instead of focusing on local details. Moreover, these tasks provide specific disease diagnoses rather than symptoms that require clinical confirmation.
Activity Detection
task (T16) targets evaluating the frequency of a specific activity or pattern. T16 evaluates the bowel function by counting the frequency of bowel sounds within a recording.
6.2 Baselines
OPERA-CT (Zhang et al., 2024)
is a contrastive learning-based pre-trained model for respiratory audio feature extraction. As the first respiratory acoustic foundation models, OPERA family models leverage quantities of breathing and cough sounds to conduct pertaining and demonstrate superior performance and generalizability against existing acoustic models pre-trained with general audio. We select OPERA-CT, the most powerful with the largest number of parameters as the baseline.
AudioMAE (Huang et al., 2022)
is a self-supervised pre-trained model for audio understanding, inspired by the Masked Autoencoders (MAE) (He et al., 2022) approach used in vision tasks. It works by randomly masking portions of the audio signal during pre-training and learning to reconstruct the missing parts. This method allows the model to capture meaningful audio features without requiring labeled data. AudioMAE has shown strong performance across various audio classification and sound event detection tasks.
CLAP (Elizalde et al., 2023)
is a language-supervised pre-trained model that bridges the gap between audio and natural language through contrastive learning. By training on paired audio and text data, CLAP learns to align audio features with corresponding textual descriptions, enabling cross-modal retrieval and classification. This approach enhances the model’s capability to interpret and relate audio content with high-level semantic information, making it a robust baseline for audio-linguistic tasks.
PANN (Kong et al., 2020)
is a convolutional neural network designed for audio pattern recognition, pre-trained on large-scale datasets for audio tagging and sound event detection. It achieves high accuracy in detecting environmental sounds, musical instruments, and other acoustic events. Its architecture is specifically tailored for efficient feature extraction from audio signals, making it a strong baseline for various audio analysis tasks. More comparisons about the number of parameters, feature dimensions, the number of overall samples, and total durations of recordings in the pretraining datasets are detailed in Table 3.
OPERA-CT AudioMAE CLAP PANN AuscultaBase-Model
@@ 6.2 Baselines
Table 3: Comparison of existing acoustic pre-trained models and AuscultaBase-Model.
6.3 Benchmarking Protocol
@@
All tasks are evaluated using the standard linear probe protocol, which freezes the pre-trained audio encoder and only trains a single fully connected layer on top of the encoded representations. Such a protocol focuses on the quality and generalizability of the encoded features and avoids the overfitting problem with a small dataset. It is noted that we adopt different data pre-processing strategies (e.g. resampling and feature extraction) for different baselines strictly following their own official implementations. As for our foundation model, the same audio pre-processing procedures are applied as in the pertaining stage. Due to the significant variation in audio lengths (e.g. 5.3$\sim$122.0 s for Cinc 2016), we split all the training set audio into equal-length segments for training. We finetune the linear head for 64 epochs with a decayed learning rate of 1e-4 by default.
Function ID Sound Type Task Type OPERA-CT AudioMAE CLAP PANN AuscultaBase-Model
Abnormality Detection T1 L MC 45.41$\pm$1.95 38.19$\pm$1.19 48.05$\pm$0.43 26.01$\pm$0.06 48.36$\pm$0.16
T2 L MC 26.20$\pm$1.09 19.77$\pm$3.01 36.42$\pm$0.10 15.54$\pm$0.02 30.28$\pm$3.11
T3 L BC 46.51$\pm$1.38 55.58$\pm$0.05 55.07$\pm$0.58 40.99$\pm$0.00 56.00$\pm$1.63
T4 L MC 33.60$\pm$0.26 28.15$\pm$1.15 33.01$\pm$0.21 17.04$\pm$0.32 33.09$\pm$1.76
T5 L MC 22.76$\pm$0.09 19.08$\pm$1.10 22.41$\pm$0.55 11.19$\pm$0.14 23.45$\pm$0.46
T6 L MC 29.63$\pm$1.39 24.66$\pm$3.47 37.61$\pm$0.30 34.92$\pm$0.01 29.61$\pm$0.19
T7 L MC 36.14$\pm$1.09 27.38$\pm$6.85 34.39$\pm$5.37 11.07$\pm$0.14 42.61$\pm$1.87
T8 H MC 44.01$\pm$1.39 41.69$\pm$2.29 47.49$\pm$4.16 28.71$\pm$0.00 39.93$\pm$0.16
Disease Diagnosis T9 L MC 22.30$\pm$3.23 21.04$\pm$1.91 35.79$\pm$9.32 24.79$\pm$3.40 31.16$\pm$0.10
T10 L ML 20.94$\pm$0.33 20.45$\pm$0.00 20.30$\pm$0.58 20.45$\pm$0.00 20.97$\pm$0.21
T11 L MC 31.46$\pm$5.22 11.83$\pm$0.00 31.69$\pm$4.05 31.45$\pm$6.85 41.54$\pm$3.44
T12 H MC 77.93$\pm$1.38 55.35$\pm$6.74 86.21$\pm$2.05 87.52$\pm$2.48 88.71$\pm$0.32
T13 H BC 66.99$\pm$1.55 73.72$\pm$1.75 64.87$\pm$1.34 65.56$\pm$0.60 76.38$\pm$1.24
T14 H BC 60.44$\pm$1.74 57.42$\pm$4.26 60.35$\pm$0.35 48.17$\pm$0.23 62.10$\pm$0.82
T15 H BC 51.02$\pm$3.75 47.27$\pm$0.00 51.02$\pm$3.75 47.27$\pm$0.00 64.96$\pm$2.22
Table 4: Macro-F1 scores for abnormality detection (T1-T8) and disease diagnosis (T9-T15) tasks. The best model for each task is in bold. We report mean and standard deviation with 3 runs.
Function ID Sound Type Task Type OPERA-CT AudioMAE CLAP PANN AuscultaBase-Model
Abnormality Detection T1 L MC 69.41$\pm$1.37 64.18$\pm$0.75 72.59$\pm$0.69 63.65$\pm$0.03 72.71$\pm$0.25
T2 L MC 64.80$\pm$0.31 63.77$\pm$0.16 65.73$\pm$0.56 63.61$\pm$0.00 66.11$\pm$0.69
T3 L BC 70.18$\pm$0.28 68.98$\pm$0.46 67.06$\pm$2.62 69.45$\pm$0.00 71.15$\pm$0.04
T4 L MC 53.08$\pm$0.20 49.62$\pm$0.12 48.93$\pm$0.29 49.61$\pm$0.00 51.91$\pm$0.31
T5 L MC 51.66$\pm$0.52 49.47$\pm$0.21 48.85$\pm$0.02 49.63$\pm$0.02 51.77$\pm$0.16
T6 L MC 47.31$\pm$1.67 42.41$\pm$16.04 56.22$\pm$0.92 58.49$\pm$0.94 51.78$\pm$5.15
T7 L MC 50.00$\pm$4.54 40.91$\pm$10.01 46.97$\pm$7.58 24.24$\pm$0.00 59.10$\pm$1.52
T8 H MC 77.53$\pm$0.32 79.75$\pm$0.63 81.49$\pm$0.48 75.63$\pm$0.00 78.49$\pm$0.31
Disease Diagnosis T9 L MC 91.33$\pm$0.27 91.60$\pm$0.00 87.27$\pm$5.64 91.50$\pm$0.37 91.60$\pm$0.26
T10 L ML 22.41$\pm$0.50 23.32$\pm$1.10 22.63$\pm$1.58 22.80$\pm$0.58 22.88$\pm$0.88
T11 L MC 38.00$\pm$2.00 42.00$\pm$0.00 45.00$\pm$1.00 43.00$\pm$3.00 50.00$\pm$2.00
T12 H MC 77.50$\pm$1.50 56.00$\pm$8.00 86.50$\pm$2.50 87.50$\pm$2.50 88.50$\pm$0.50
T13 H BC 83.23$\pm$0.31 81.99$\pm$0.62 82.14$\pm$0.16 78.11$\pm$1.09 86.50$\pm$0.47
T14 H BC 61.39$\pm$0.95 59.65$\pm$5.22 60.76$\pm$0.63 55.70$\pm$0.00 63.29$\pm$0.63
T15 H BC 89.87$\pm$0.21 89.66$\pm$0.00 89.87$\pm$0.22 89.66$\pm$0.00 90.52$\pm$0.00
Table 5: Micro-F1 scores for abnormality detection (T1-T8) and disease diagnosis tasks (T9-T15). The best model is in bold, and the mean and standard deviation are reported with 3 runs.
6.4 Metric
Figure 2: Accuracy scores for the activity detection task. The best model is in bold, and the error bars are reported with 3 runs.

### v2

@@ 4.1 AuscultaCorpus: Pretraining Datasets
Table 2: Statistics of the AuscultaCorpus (SR: sampling rate; Duration: mean [min, max]), where the sound type is categorized into lung (L), heart (H), and bowel (B) sounds.
Given the heterogeneity in auscultation positions, sampling rates, and recording durations of those datasets, each recording is first resampled to 16kHz and converted to a mono channel. Subsequent processing involves trimming leading and trailing silences to enhance recording quality. The refined audio segments are then transformed into 64-bin LogMel spectrograms using a Hanning window with a 32ms stride. Finally, Min–Max Normalization is applied to standardize the spectrograms prior to their use in pretraining the auscultation foundation model.
4.2 AuscultaBase: Auscultation Foundation Model
4.2.1 Training Algorithm
Given the inherent similarities among body sound recordings, distinguishing subtle differences is crucial for detecting abnormalities. To this end, we employ a contrastive learning framework during pretraining. Specifically, two segments extracted from the same recording are designated as positive pairs, whereas segments drawn from different recordings serve as negative pairs. Given two segments $x$ and $x^{\prime}$, the features will first be extracted with an encoder $f(\cdot)$ from them, and then a projector $g(\cdot)$ will map the features into a lower dimensional space to obtain the final representation. To improve the quality and discriminative power of the learned representations, an extra linear layer $W$ is added on top of the whole model for the positive/negative anchor $x^{\prime}$. Then the bilinear similarity $s(\cdot)$ is calculated as
$s(x,x^{\prime})=g(f(x))^{T}Wg(f(x^{\prime})).$ (1)
This approach compels the encoder to discern fine-grained acoustic variations, thereby yielding robust and generalizable features suitable for subsequent downstream analyses.
4.2.2 Model Architecture
In view of the variable lengths of body sound recordings, we adopt a transformer-based network, following the design principles outlined in Chen et al. (2022), to serve as our encoder. The extracted features are subsequently projected into a lower-dimensional space via a dedicated projection layer. Within this space, a bilinear similarity metric is computed between representations, and the optimization objective is formulated to enhance the similarity of positive pairs while reducing that of negative pairs. To manage the inherent variability in audio durations within each batch, we implement random cropping to obtain standardized training segments. Furthermore, to ensure a sufficiently challenging optimization process, separate dataloaders are established for each dataset so that all audio clips within a given batch originate from the same source; batches are then sampled randomly across these dataloaders.
4.2.3 Training Details
We pre-train our model using AuscultaCorpus, a combination of the 11 datasets. We apply validation with the validation sets of the datasets (e.g. SPRSound, HF Lung, ICBHI 2017, and Bowel Sound). For those datasets that do not have an existing validation set, we randomly split 10% of the data to serve for validation. To adapt to the varied lengths of recordings, we adopt random cropping for each audio and feed them to the pertaining. according to the minimum duration of the recordings for each dataset, different crop lengths are set. Specifically, the cropped length for Korean and Bowel Sound datasets is set as 320 ms, while it is 640 ms for others. The encoded representations are projected into a 512-dimensional space, and the model is trained over 200 epochs using the Adam optimizer. The initial learning rate is 1e-04 and decays by a factor of 0.99 at each epoch. Additionally, as a usual augmentation technique, we adjust the loudness of the recording (Schlüter & Grill, 2015) with a random scale factor ranging from 0.9 to 1.1. Moreover, SpecAugment (Park et al., 2019) is also adopted.
4.3 AuscultaBench: Auscultation Benchmark
@@
ID Function Dataset Sound Type Task Type Class Name Data Distribution
T1 AD SPRSound L MC Normal / Adventitious / Poor Quality 2324 / 1000 / 230
T2 AD SPRSound L MC Normal / CAS / DAS / CAS & DAS / Poor Quality 2324 / 368 / 480 / 152 / 230
T3 AD HF Lung L BC Normal / Abnormal 52444 / 29489
T4 AD HF Lung L MC Inhalation / Exhalation / CAS / DAS 34095 / 18349 / 13883 / 15606
T5 AD HF Lung L MC Inhalation / Exhalation / Wheeze / Stridor / Rhonchi / Crackle 34095 / 18349 / 8457 / 686 / 4740 / 15606
T6 AD ICBHI 2017 L MC Normal / Crackle / Wheeze / Crackle&Wheeze 3642 / 1864 / 886 / 506
T7 AD Lung Sound L MC Normal / Crepitation / Wheeze / Crackle / Bronchi / Wheeze & Crackle / Bronchi&Crackle 105 / 69 / 123 / 24 / 3 / 6 / 6
T8 AD Circor 2022 H MC Murmur Present / Absent / Unknown 363 / 2391 / 156
T9 AD Bowel Sound B BC Bowel sound Present / Absent 1283 / 323
T10 DD ICBHI 2017 L MC Healthy / Bronchiectasis / Bronchiolitis / COPD / Pneumonia / URTI 35 / 16 / 13 / 793 / 37 / 23
@@
T11 DD Lung Sound L ML Normal / Asthma / Pneumonia / COPD / BRON / Heart failure / Lung fibrosis / Pleural effusion 105 / 99 / 15 / 33 / 9 / 63 / 18 / 6
T12 DD Respiratory Database@TR L MC COPD0 / COPD1 / COPD2 / COPD3 / COPD4 72 / 60 / 84 / 84 / 204
@@
T13 DD Korean H MC Normal / Aortic Stenosis / Mitral Regurgitation / Mitral Stenosis / Murmur in Systole 200 / 200 / 200 / 200 / 200
T14 DD Cinc 2016 H BC Normal / Abnormal 2575 / 665
T15 DD Circor 2022 H BC Normal / Abnormal 1632 / 1531
T16 DD HSDReport H BC Normal / Abnormal 247 / 2028
@@
Table 3: Downstream task characteristics of AuscultaBench grouped into abnormality detection (AD: T1-T9) and disease diagnosis tasks(DD: T10-T16). The task type is categorized into binary classification (BC), multi-class classification (MC), and multi-label classification (ML). Data Distribution denotes the number of samples corresponding to the class names.
4.3.1 Tasks
@@
Our benchmark comprises 16 sub-tasks encompassing binary classification (BC), multi-class classification (MC), and multi-label classification (ML), as detailed in Figure 2. Please refer to Table 3 for more details. These tasks are organized by their functional objectives into abnormality detection and disease diagnosis.
@@
tasks (T1-T9) are designed to discern whether a body sound is normal and, if abnormal, to specify the type of abnormality or normality present. For example, task T8 is aimed at identifying the presence of murmurs in heart sounds, whereas task T4 requires the model to determine whether the patient is inhaling or exhaling and, if adventitious sounds are detected, to further classify them as continuous (CAS) or discrete (DAS). While abnormalities in body sounds may signal potential health concerns, these tasks do not substitute for a definitive clinical diagnosis, which must be confirmed by a clinician.
@@
tasks (T10-T16) focus on determining whether a patient is healthy or affected by a specific disease. For instance, task T13 distinguishes between healthy heart sounds and those indicative of cardiac disease, and task T12 involves identifying the stage of chronic obstructive pulmonary disease (COPD), ranging from 0 to 4. Notably, we have excluded the “LRTI” and “Asthma” categories from the original dataset for task T10 due to their minimal representation, which precludes their appearance in the test set. These tasks require the model to analyze the entire recording to detect potential disease signatures rather than relying solely on localized features, and they provide disease-specific diagnoses rather than preliminary symptomatic observations.
4.3.2 Data Pre-processing
we adopt the validation set of the pretraining stage as the test set, which is always held out for pretraining and downstream fine-tuning. Considering the varied lengths of different recordings even within the same dataset, each audio in the training set is chunked and padded into the same length. Specifically, each recording of all the datasets is chunked into 8s to satisfy the input requirements of different models except for T3-T5 (HF Lung dataset), T9 (Bowel Sound dataset), T13 (Korean dataset), and T16 (HSDReport dataset). As for the HF Lung, Bowel Sound, Korean, and HSDReport datasets, the maximum durations are set as 2s, 4s, 75s, and 2s respectively because their recordings are rather short or long as demonstrated in Table 2.
OPERA-CT AudioMAE CLAP PANN AuscultaBase
@@ 4.3.2 Data Pre-processing
Table 4: Comparison of existing acoustic pre-trained models and AuscultaBase.
4.3.3 Baselines
OPERA-CT
(Zhang et al., 2024) is a contrastive learning-based pre-trained model designed specifically for respiratory audio feature extraction. As one of the first respiratory acoustic foundation models, the OPERA family leverages extensive collections of breathing and cough sounds to pretrain its architecture, thereby demonstrating superior performance and generalizability compared to conventional acoustic models pre-trained on general audio. In this study, we adopt OPERA-CT, the variant with the largest number of parameters, as a baseline.
AudioMAE
(Huang et al., 2022) is a self-supervised pre-trained model for audio understanding, inspired by the Masked Autoencoders (MAE) approach originally developed for vision tasks (He et al., 2022). It operates by randomly masking segments of the audio signal during pretraining and then learning to reconstruct the missing portions. This strategy enables the model to capture salient audio features without reliance on labeled data, and it has demonstrated robust performance across a range of audio classification and sound event detection tasks.
CLAP
(Elizalde et al., 2023) is a language-supervised pre-trained model that bridges audio and natural language via contrastive learning. By training on paired audio and text data, CLAP aligns audio features with their corresponding textual descriptions, facilitating cross-modal retrieval and classification. This approach enhances the model’s ability to interpret and relate audio content to high-level semantic information, making it a strong baseline for audio-linguistic tasks.
PANN
(Kong et al., 2020) is a convolutional neural network tailored for audio pattern recognition and pre-trained on large-scale datasets for audio tagging and sound event detection. It achieves high accuracy in detecting environmental sounds, musical instruments, and other acoustic events due to its efficient architecture for audio feature extraction. Further comparisons concerning the number of parameters, feature dimensions, total sample counts, and the aggregate durations of recordings in the pretraining datasets are detailed in Table 4.
4.3.4 Benchmarking Protocol
@@
All tasks are evaluated using the standard linear probe protocol, which freezes the pre-trained audio encoder and only trains a single fully connected layer on top of the encoded representations. Such a protocol focuses on the quality and generalizability of the encoded features and avoids the overfitting problem with a small dataset. It is noted that we adopt different data pre-processing strategies (e.g. resampling and feature extraction) for different baselines strictly following their own official implementations. As for our foundation model, the same audio pre-processing procedures are applied as in the pertaining stage. Due to the significant variation in audio lengths (e.g. 5.3$\sim$122.0 seconds for Cinc 2016), the training set recordings are segmented into equal-length portions prior to training. The linear head is then fine-tuned for 64 epochs using a decaying learning rate, with an initial value of 1e-4.
4.3.5 Evaluation Metric